OpenStudy (anonymous):

solve 6x^2 + 6x -12 = 0

OpenStudy (campbell_st):

start by dividing each term by 6 and you have \[x^2 + x - 2 = 0\] now you can solve by factoring... find the factors of -2 that add to 1...the larger factor is positive and the smaller is negative hope this helps

OpenStudy (anonymous):

so x = 2 and x = -1 ?

OpenStudy (campbell_st):

well there are the factors so you have (x +2)(x -1) = 0 now you need to find the solutions by solving x + 2 = 0 and x - 1 = 0

OpenStudy (anonymous):

understood. thanks